### v1

# An uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero

Abstract.
We construct an uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero (and hence of Lebesgue measure zero).
We denote by $\mathbb{N}_{0}$ the set of non-negative integers, by $\mathbb{N}$ the set of positive integers and by $\mathbb{Z}$ the set of integers. Let
$S:=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}=\{0,1,2,2^{2},2^{3},\ldots\}.$
For every $n\in\mathbb{N}$ by $nS$, we denote the $n^{\text{th}}$ sumset of $S$, i.e.,
$nS:=\{s_{1}+\cdots+s_{n}:s_{1},\ldots,s_{n}\in S\}.$
For every $n\in\mathbb{N}$ we define
$A_{n}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}:(x_{k})_{k\in nS}\text{is a bounded sequence of integers}\right\}.$
Theorem 1.
Let $n\in\mathbb{N}$. The set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$.
Proof.
We first show that $A_{n}$ is a subgroup of $\mathbb{R}$. Trivially, $0\in A_{n}$ and $-x\in A_{n}$ whenever $x\in A_{n}$. Thus, it suffices to show that $A_{n}$ is closed under addition. Let $x,y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}$
where $(x_{k})_{k\in nS}$ and $(y_{k})_{k\in nS}$ are bounded sequences of integers. It follows that
$x+y=\sum\limits_{k\in nS}\frac{x_{k}+y_{k}}{2^{k}}$
is also an element of $A_{n}$.
Clearly, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of $1$’s, the set of real numbers of the form
$\sum\limits_{k\in nS}\frac{z_{k}}{2^{k}}\quad\mbox{with }z_{k}\in\{0,1\}$
is an uncountable subset of $A_{n}$. ∎
We now proceed to show that the set $A_{n}$ has Hausdorff dimension zero for every $n\in\mathbb{N}$. For every $n\in\mathbb{N}$ write
$A_{n}:=\bigcup_{t=1}^{\infty}A_{n,t}$
where for every $t\in\mathbb{N}$ we let
$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:x_{k}\in\mathbb{Z}\mbox{ and }|x_{k}|\leq t\mbox{ for all }k\in nS\right\}.$
We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
Lemma 1.
For $n\in\mathbb{N}$ and $z\geq 1$, define
$g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}.$
Then we have the bound
$g_{n}(z)\leq(2+\log_{2}z)^{n}.$
Proof.
Clearly, for every $z\geq 1$, we have
$g_{n}(z)\leq g_{1}(z)^{n}\mbox{ and }g_{1}(z)\leq 2+\text{log}_{2}z.$
The result follows. ∎
Theorem 2.
The set $A_{n,t}$ has Hausdorff dimension zero for all $n,t\in\mathbb{N}$.
Proof.
Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Let $x\in nS$ and write
$x=\sum\limits_{i=1}^{\infty}\frac{a_{i}}{2^{k_{i}}}$
with $a_{i}\in\mathbb{Z}$ and $|a_{i}|\leq t$. Pick $j\in\mathbb{N}$ and split $x$ into
$x=m_{x}+e_{x},$
where
$m_{x}:=\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}\quad\mbox{and}\quad e_{x}:=\sum\limits_{i=j}^{\infty}\frac{a_{i}}{2^{k_{i}}}.$
Then using the triangle inequality, we have the bound
$|e_{x}|\leq t\sum\limits_{i=j}^{\infty}\frac{1}{2^{k_{i}}}\leq t\sum\limits_{h=k_{j}}^{\infty}\frac{1}{2^{h}}=\frac{t}{2^{k_{j}-1}}.$
Let $r_{j}=t/2^{k_{j}-1}$. It follows that
$A_{n,t}\subseteq\bigcup_{m\in X}I_{m},$
where
$X:=\left\{\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}:a_{i}\in\mathbb{Z},\ |a_{i}|\leq t\right\}$
and
$I_{m}:=[m-r_{j},m+r_{j}].$
Then #$\emph{{\#}}(X)\leq(2t+1)^{j-1}$ and for every $m\in X$ we see that $\mu(I_{m})=2r_{j}$ where $\mu(I_{m})$ is the length of the interval $I_{m}$.
Let $d>0$ be a real number, let $\delta_{j}:=4r_{j}$, and let $\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})$ be defined as in [2]. We see that
$\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})\leq\sum\limits_{m\in X}\mu(I_{m})^{d}=\emph{{\#}}(X)(2r_{j})^{d}\leq\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}.$
We claim that
(1) $\lim\limits_{j\rightarrow\infty}\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}=0.$
Since $\delta_{j}\to 0$ as $j\to\infty$, this implies that $\mathcal{H}^{d}(A_{n,t})=0$ where $\mathcal{H}^{d}(A_{n,t})$ is the $d$-dimensional Hausdorff measure of $A_{n,t}$. (For the definition of this measure and related terms, see [2].) Since this holds for all $d>0$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
Equation (1) above is seen as follows. First, it is convenient to note that (1) holds if
$\lim\limits_{j\rightarrow\infty}\frac{(3t)^{j}}{2^{k_{j}d}}=0.$
By taking logarithms, we see that (1) holds if
$\lim\limits_{j\rightarrow\infty}\big{(}k_{j}d-j\log_{2}(3t)\big{)}=\infty.$
As $d>0$, we let $c$ be the constant $(1/d)\log_{2}(3t)$ and we see that (1) holds if
(2) $\lim\limits_{j\to\infty}(k_{j}-cj)=\infty.$
Now using Lemma 1, we have
$j=g_{n}(k_{j})\leq(2+\log_{2}k_{j})^{n},$
which implies
$k_{j}\geq 2^{\sqrt[n]{j}-2}.$
Hence, (2) follows, which justifies our claim. ∎
Let $n\in\mathbb{N}$. From Theorems 1 and 2 it follows that the set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. As $0\in S$, we have $nS\subseteq(n+1)S$, and hence $A_{n}\subseteq A_{n+1}$. Let
$A:=\bigcup_{n=1}^{\infty}A_{n}.$
We deduce that the set $A$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that $A$ is in fact a subring. To this end, we need the following result.
Lemma 2.
Let $n\in\mathbb{N}$ and let $a_{n}(s)$ be the number of representations of an element $s\in\mathbb{N}$ as a sum $s=t_{1}+t_{2}+\cdots+t_{n}$ of $n$ elements $t_{1},t_{2},\ldots,t_{n}\in S$. Then there exists $c_{n}\in\mathbb{R}$ such that $a_{n}(s)\leq c_{n}$ for every $s\in\mathbb{N}$.
Proof.
Clearly $a_{1}(s)\leq 1$ for every $s\in\mathbb{N}$ and hence, we may take $c_{1}:=1$. Now assume that $n\geq 2$ and that there exists $c_{n-1}\in\mathbb{R}$ such that $a_{n-1}(s)\leq c_{n-1}$ for every $s\in\mathbb{N}$. Let $s\in\mathbb{N}$ and suppose
$s=t_{1}+\cdots+t_{n}$
is a representation of $s$ as a sum of $n$ elements of $S$. Then for at least one $i\in\{1,\ldots,n\}$, we have $t_{i}\geq s/n$ and hence $t_{i}\neq 0$. So there exists $k\in\mathbb{N}_{0}$ such that $t_{i}=2^{k}$. As $s/n\leq t_{i}\leq s$, we observe that $\log_{2}s-\log_{2}n\leq k\leq\log_{2}s$. So if
$B_{n,s}:=[\,\log_{2}s-\log_{2}n,\,\log_{2}s\,]\cap\mathbb{N}_{0},$
then we see that
$a_{n}(s)\leq n\sum\limits_{k\in B_{n,s}}a_{n-1}(s-2^{k})\leq n(1+\log_{2}n)c_{n-1}.$
So we are done by taking $c_{n}:=n(1+\log_{2}n)c_{n-1}$. ∎
Theorem 3.
The set $A$ is a a subring of $\mathbb{R}$.
Proof.
It suffices to show that $1\in A$ and that $A$ is closed under multiplication. We see that
$1=\frac{1}{2^{0}}+\sum\limits_{k\in S\setminus\{0\}}\frac{0}{2^{k}}\in A_{1}.$
To establish that $A$ is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exists $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}$
where $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$ are bounded sequences of integers. Now
$xy=\left(\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\right)\left(\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}\right)=\sum\limits_{(k,\ell)\in mS\times nS}\frac{x_{k}y_{\ell}}{2^{k+\ell}}=\sum\limits_{r\in(m+n)S}\frac{z_{r}}{2^{r}}$
where
$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=t_{1}+t_{2}+\cdots+t_{m+n}$ where $t_{1},t_{2},\ldots,t_{m+n}\in S$. Hence, we deduce that the sequence $(z_{r})_{r\in(m+n)S}$ is bounded as a consequence of Lemma 2 and the boundedness of the sequences $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$. It follows that $xy\in A_{m+n}$. This establishes that $A$ is closed under multiplication. ∎
We now proceed to show that the only rational numbers in $A$ are the dyadic rationals. In particular, it will follow that the ring $A$ is not a field.
Lemma 3.
Let $r,n\in\mathbb{N}$. Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Then there exists $m\in\mathbb{N}$ such that $k_{m+1}-k_{m}>r$.
Proof.
For every $z\in\mathbb{N}$, let $g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}$. If the conclusion is false, then $g_{n}(z)\geq z/r$ for every $z\in\mathbb{N}$. This contradicts Lemma 1. ∎
Proposition 1.
Let $x\in A$ and $l\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $l$.
Proof.
Let $x\in A_{n,t}$. Then there exists a sequence of integers $(x_{k})_{k\in nS}$ with $|x_{k}|\leq t$ for all $k\in nS$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}=\sum\limits_{i=1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}}}.$
Let $m\in\mathbb{N}$. We observe that
$2^{k_{m}}x=q_{m}+r_{m}$
where
$q_{m}:=\sum\limits_{i=0}^{m}2^{k_{m}-k_{i}}x_{k_{i}}\mbox{ and }r_{m}:=\sum\limits_{i=m+1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}-k_{m}}}.$
Then $q_{m}\in\mathbb{Z}$ and we have
$\left|r_{m}\right|\leq~{}t\sum\limits_{j=k_{m+1}-k_{m}}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{k_{m+1}-k_{m}}}.$
By Lemma 3, there exists $m\in\mathbb{N}$ such that $2^{k_{m+1}-k_{m}}>2^{l+1}t$ and so $|r_{m}|<1/2^{l}$. Hence,
$\text{frac}\left(2^{k_{m}}x\right)=\begin{cases}r_{m},&\mbox{ if }r_{m}\in\left[\,0,\frac{1}{2^{l}}\right);\\ 1+r_{m},&\mbox{ if }r_{m}\in\left(-\frac{1}{2^{l}},0\right);\end{cases}$
and so
$\text{frac}(2^{k_{m}}x)\in\left[0,\tfrac{1}{2^{l}}\right)\cup\left(1-\tfrac{1}{2^{l}},1\right),$
where frac$(z):=z-[z]$ is the fractional part of $z\in\mathbb{R}$. It follows that the digits after the dot in a binary expansion of $x$ from the $(k_{m}+1)^{\text{th}}$ position to the $(k_{m}+l)^{\text{th}}$ position are all equal. ∎
Corollary 1.
A number $x\in A$ is rational if and only if there exists $a\in\mathbb{Z}$ and $k\in\mathbb{N}_{0}$ such that $x=a/2^{k}$.
Proof.
From Proposition 1, we see that the ring $A$ only contains those rational numbers which have a finite binary expansion. The result follows. ∎
Remark 1.
The above construction of an uncountable subring of $\mathbb{R}$ of Hausdorff dimension zero goes through if the initial set $S=\{0,1,2,2^{2},2^{3},\ldots\}$ is replaced by a set $T\cup\{0\}$ where $T$ is a subset of $\mathbb{N}$ which has the following property: There exists $b\in\mathbb{R}$ such that for every $n\in\mathbb{N}$, the cardinality of the set $\{\log t:t\in T\}\cap[n,n+1)$ is less than $b$.
Acknowledgment: We thank Prof. Shrikrishna G. Dani for sharing with us a preprint of [1] in which we see a different construction of an uncountable subgroup of $\mathbb{R}$ having Hausdorff dimension zero.