### v2

# An uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero

Abstract.
We construct a subring as mentioned in the title (hence this subring has Lebesgue measure zero).
1. Preliminaries
For a subset $B\subseteq\mathbb{R}$ we let $d(B)$ denote the diameter of $B$. Given a countable collection of subsets $\mathcal{C}\subseteq\mathcal{P}(\mathbb{R})$ and $s\geq 0$ we define
$d(s,\mathcal{C}):=\sum_{B\in\mathcal{C}}d(B)^{s}.$
We note that $d(s,\mathcal{C})\in[0,\infty]$ and does not depend on the ordering of the subsets in the collection $\mathcal{C}$. For $\delta>0$, we say that a collection of subsets $\mathcal{C}$ is a $\delta$-cover of a subset $F\subseteq\mathbb{R}$ if for every $B\in\mathcal{C}$ we have $d(B)\leq\delta$ and if $F\subseteq\cup\,\mathcal{C}$.
Given a subset $F\subseteq\mathbb{R}$ and $s,\delta\in\mathbb{R}$ with $s\geq 0$, and $\delta>0$, we define
$\mathcal{H}^{s}_{\delta}(F)=\inf\big{\{}d(s,\mathcal{C}):\mathcal{C}\mbox{ is a countable $\delta$-cover of $F$}\big{\}}.$
If $\delta_{1}<\delta_{2}$, then we see that $\mathcal{H}^{s}_{\delta_{1}}(F)\geq\mathcal{H}^{s}_{\delta_{2}}(F)$. The $s$-dimensional Hausdorff measure $\mathcal{H}^{s}(F)$ is defined to be the limit of $\mathcal{H}_{\delta}^{s}(F)$ as $\delta\to 0^{+}$.
It can be shown (see [4]) that there exists $\alpha\in[0,1]$ such that for every $s\in[0,\alpha)$ we have $\mathcal{H}^{s}(F)=\infty$, and for every $s>\alpha$ we have $\mathcal{H}^{s}(F)=0$. The number $\alpha$ is called the Hausdorff dimension of $F$ and is denoted by $\dim_{\mathcal{H}}F$. If $F=\cup\,\mathcal{C}$ where $\mathcal{C}$ is a countable collection of subsets of $\mathbb{R}$, then $\dim_{\mathcal{H}}F=\sup\big{\{}\dim_{\mathcal{H}}A:A\in\mathcal{C}\big{\}}$.
2. Construction
We give an explicit construction of a proper subring of $\mathbb{R}$ which is a Borel set. We also show explicitly that it has Hausdorff dimension zero.
We let $\mathbb{N}_{0}$ denote the set of non-negative integers, $\mathbb{N}$ denote the set of positive integers, and $\mathbb{Z}$ denote the set of integers. For every $a,b\in\mathbb{Z}$ we let $a,b]_{\mathbb{Z}}=\{k\in\mathbb{Z}:a\leq k\leq b\}$. Let
$S:=\{0,1,2,2^{2},2^{3},\ldots\}=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}.$
For every $n\in\mathbb{N}$ we let the $n^{\text{th}}$ sumset of $S$ be defined as
$nS:=\{s_{1}+\cdots+s_{n}:s_{1},\ldots,s_{n}\in S\}.$
For every $n\in\mathbb{N}$ we define
[…]
Theorem 1.
Let $n\in\mathbb{N}$. The set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$.
Proof.
We first show that $A_{n}$ is a subgroup of $\mathbb{R}$. Trivially, we have $0\in A_{n}$, and that $x\in A_{n}$ whenever $x\in A_{n}$. Thus, it suffices to show that $A_{n}$ is closed under addition. Let $x,y\in A_{n}$. Therefore, there exists $t\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k},y_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}.$
It follows that
$x+y=\sum\limits_{k\in nS}\frac{x_{k}+y_{k}}{2^{k}}$
is also an element of $A_{n}$ since $x_{k}+y_{k}\in[-2t,2t\,]_{\mathbb{Z}}$ for every $k\in nS$. Also, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of ones, the set of real numbers of the form
[…]
is an uncountable subset of $A_{n}$. ∎
We now proceed to show that the set $A_{n}$ has Hausdorff dimension zero for every $n\in\mathbb{N}$. For every $n\in\mathbb{N}$ write
$A_{n}:=\bigcup_{t=1}^{\infty}A_{n,t}$
where for every $t\in\mathbb{N}$ we let
$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:\mbox{ for every }k\in nS\mbox{ we have }x_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}.$
For every $n,\ell\in\mathbb{N}$ we let $(nS)_{\leq\ell}=\{k\in nS:k\leq\ell\}$ and we let $(nS)_{>\ell}=\{k\in nS:k>\ell\}$. We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
Lemma 1.
For every $n,\ell\in\mathbb{N}$ we define
$g_{n}(\ell):=\emph{{\#}}(nS)_{\leq\ell}.$
Then we have the bound
$g_{n}(\ell)\leq(2+\log_{2}\ell)^{n}.$
Proof.
We observe that for every $z\in\mathbb{N}$, we have
$g_{n}(\ell)\leq g_{1}(\ell)^{n}\mbox{ and }g_{1}(\ell)\leq 2+\log_{2}\ell.$
The result follows. ∎
Theorem 2.
The set $A_{n,t}$ has Hausdorff dimension zero for all $n,t\in\mathbb{N}$.
Proof.
Let $x\in A_{n,t}$. Then for every $k\in nS$ there exist $a_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{a_{k}}{2^{k}}.$
Pick $\ell\in\mathbb{N}$ and write $x$ as
$x=y+z,$
where
$y:=\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}\quad\mbox{and}\quad z:=\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{a_{k}}{2^{k}}.$
Then using the triangle inequality, we have the bound
$|z|\leq t\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{1}{2^{k}}\leq t\sum\limits_{k=\ell+1}^{\infty}\frac{1}{2^{k}}=\frac{t}{2^{\ell}}.$
Let
$B:=\left\{\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}:\mbox{ for every }k\in(nS)_{\leq\ell}\mbox{ we have }a_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}$
and let $r=t/2^{\ell}$. It follows that
$A_{n,t}\subseteq\bigcup_{b\in B}I_{b}\mbox{ where }I_{b}=[b-r,b+r]\mbox{ for every }b\in B.$
Let $c_{\ell}=\emph{{\#}}\,(nS)_{\leq\ell}$. Then $\emph{{\#}}B\leq(2t+1)^{c_{\ell}}$ and for every $b\in B$ we see that $\mu(I_{b})=2r$ where $\mu$ is the Lebesgue measure on $\mathbb{R}$. Let $s\in(0,1]$. We see that
$\mathcal{H}_{2r_{\ell}}^{s}(A_{n,t})\leq\sum\limits_{b\in B}\mu(I_{b})^{s}=\emph{{\#}}B\,(2r)^{s}\leq(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}.$
We claim that
(1) $\lim\limits_{\ell\rightarrow\infty}(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}=0.$
Since $r\to 0$ as $\ell\to\infty$, this implies that $\mathcal{H}^{s}(A_{n,t})=0$. As this holds for all $s\in(0,1]$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
Equation (1) above is seen as follows. First, it is convenient to note that (1) holds if
$\lim\limits_{\ell\to\infty}\frac{(3t)^{c_{\ell}}}{2^{\ell s}}=0.$
Now, by taking logarithms, we see that (1) holds if
$\lim\limits_{\ell\to\infty}\Big{(}\ell s-c_{\ell}\log_{2}(3t)\Big{)}=\infty.$
Let $\alpha=\frac{\log_{2}(3t)}{s}$. As $s>0$, we see that (1) holds if
(2) $\lim\limits_{\ell\to\infty}\Big{(}\ell-\alpha c_{\ell}\Big{)}=\infty.$
Now using Lemma 1, we have
$c_{\ell}\leq(2+\log_{2}\ell)^{n},$
which implies
$\ell\geq 2^{c_{\ell}^{r}-2}$
where $r=1/n$. Hence, it follows that (2) holds. This justifies our claim. ∎
Let $n\in\mathbb{N}$. From Theorems 1 and 2 it follows that the set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. As $0\in S$, we have $nS\subseteq(n+1)S$, and hence $A_{n}\subseteq A_{n+1}$. Let
$A:=\bigcup_{n=1}^{\infty}A_{n}.$
We deduce that the set AA is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that AA is in fact a subring. To this end, we need the following result.
Lemma 2.
Let $n\in\mathbb{N}$ and let $c_{n}(k)$ be the number of representations of an element $k\in\mathbb{N}$ as a sum $k=s_{1}+s_{2}+\cdots+s_{n}$ of $n$ elements of $S$. Then there exists $b_{n}\in\mathbb{N}$ such that $c_{n}(k)\leq b_{n}$ for every $k\in\mathbb{N}$.
Proof.
Clearly $c_{1}(k)\leq 1$ for every $k\in\mathbb{N}$ and hence, we may take $b_{1}:=1$. Now assume that $n\geq 2$ and that there exists $b_{n-1}\in\mathbb{N}$ such that $c_{n-1}(k)\leq b_{n-1}$ for every $k\in\mathbb{N}$. Let $k\in\mathbb{N}$ and suppose
$k=s_{1}+\cdots+s_{n}$
is a representation of $k$ as a sum of $n$ elements of $S$. Then for at least one $i\in\{1,\ldots,n\}$ we have $s_{i}\geq k/n$ and hence $s_{i}\neq 0$. So there exists $m\in\mathbb{N}_{0}$ such that $s_{i}=2^{m}$. As $k/n\leq s_{i}\leq k$, we observe that $\log_{2}k-\log_{2}n\leq m\leq\log_{2}k$. Let
$I_{n,k}:=[\,\log_{2}k-\log_{2}n,\,\log_{2}k\,]\cap\mathbb{N}_{0}.$
Then we see that
$c_{n}(k)\leq n\sum\limits_{m\in I_{n,k}}c_{n-1}(k-2^{m})\leq nb_{n-1}\emph{{\#}}I_{n,k}\leq nb_{n-1}(1+\log_{2}n).$
So we are done by taking $b_{n}\in\mathbb{N}$ such that $b_{n}\geq nb_{n-1}(1+\log_{2}n)$. ∎
Theorem 3.
The set AA is a a subring of $\mathbb{R}$.
Proof.
It suffices to show that $1\in A$ and that AA is closed under multiplication. We see that
$1=\frac{1}{2^{0}}+\sum\limits_{k\in S_{>0}}\frac{0}{2^{k}}\in A_{1}.$
To establish that AA is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exist $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore, there exist $t,t^{\prime}\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ and $y_{k}\in[-t^{\prime},t^{\prime}]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}.$
Now
$xy=\left(\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\right)\left(\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}\right)=\sum\limits_{(k,\ell)\in mS\times nS}\frac{x_{k}y_{\ell}}{2^{k+\ell}}=\sum\limits_{r\in(m+n)S}\frac{z_{r}}{2^{r}}$
where for every $r\in(m+n)S$ we let
$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\vskip 3.0pt plus 1.0pt minus 1.0pt\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=s_{1}+s_{2}+\cdots+s_{m+n}$ where $s_{1},s_{2},\ldots,s_{m+n}\in S$. Hence, by Lemma 2 we deduce that there exists $b\in\mathbb{N}$ such that for every $r\in(m+n)S$ we have $|z_{r}|\leq btt^{\prime}$. It follows that $xy\in A_{m+n}$. This establishes that AA is closed under multiplication. ∎
We now proceed to show that the only rational numbers in AA are the dyadic rationals. In particular, it will follow that the ring AA is not a field.
Lemma 3.
Let $b,n\in\mathbb{N}$. Then there exist two consecutive elements $k,k^{\prime}\in nS$ such that $k^{\prime}-k>b$.
Proof.
For every $\ell\in\mathbb{N}$ we let $c_{\ell}:=\emph{{\#}}\,(nS)_{\leq\ell}$. If the conclusion is false, then $c_{\ell}\geq\ell/b$ for every $\ell\in\mathbb{N}$. This contradicts Lemma 1. ∎
Proposition 1.
Let $x\in A$ and $\ell\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $\ell$.
Proof.
Let $x\in A_{n,t}$. Then for every $k\in nS$ there exists $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}.$
Let $m\in nS$. We observe that
$2^{m}x=q_{m}+r_{m}$
where
$q_{m}:=\sum\limits_{k\in(nS)_{\leq m}}2^{m-k}x_{k}\mbox{ and }r_{m}:=\sum\limits_{k\in(nS)_{>m}}\frac{x_{k}}{2^{k-m}}.$
Then $q_{m}\in\mathbb{Z}$ and we have
$|r_{m}|\leq~t\sum\limits_{j=m^{\prime}-m}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{m^{\prime}-m}}$
where $m,m^{\prime}\in nS$ are consecutive elements.
By Lemma 3, there exist consecutive elements $k,k^{\prime}\in nS$ such that $2^{k^{\prime}-k}>2^{\ell+1}t$ and so we get $|r_{k}|<\frac{1}{2^{\ell}}$. Hence,
$\text{frac}\left(2^{k}x\right)=\begin{cases}r_{k},&\mbox{ if }r_{k}\in\left[\,0,\frac{1}{2^{\ell}}\right);\\ 1+r_{k},&\mbox{ if }r_{k}\in\left(-\frac{1}{2^{\ell}},0\right);\end{cases}$
and so
$\text{frac}(2^{k}x)\in\left[0,\tfrac{1}{2^{\ell}}\right)\cup\left(1-\tfrac{1}{2^{\ell}},1\right),$
[…]
Corollary 1.
A number $x\in A$ is rational if and only if there exists $a\in\mathbb{Z}$ and $k\in\mathbb{N}_{0}$ such that $x=a/2^{k}$.
Proof.
From Proposition 1, we see that the ring AA only contains those rational numbers which have a finite binary expansion. The result follows. ∎
Remark 1.
The above construction of an uncountable subring of $\mathbb{R}$ of Hausdorff dimension zero goes through if the initial set $S=\{0,1,2,2^{2},2^{3},\ldots\}$ is replaced by a set $T\cup\{0\}$ where $T$ is a subset of $\mathbb{N}$ which has the following property: There exists $b\in\mathbb{R}$ such that for every $n\in\mathbb{N}$, the cardinality of the set $\{\log t:t\in T\}\cap[n,n+1)$ is less than $b$.
3. Concluding remarks
In 1966, Erdős and Volkmann [2] showed that for every $d\in[0,1]$ there exists a subgroup of $\mathbb{R}$ which is Borel measurable and has Hausdorff dimension $d$. (Example 12.4 of [4]). In an unpublished work (see [3], pg. 212), using the Continuum Hypothesis, R. O. Davies showed that for every $d\in[0,1]$ there exists a subring of $\mathbb{R}$ which has Hausdorff dimension $d$ and which is not a Borel set. In [1], it is shown that every proper subring of $\mathbb{R}$ which is a Borel set, has Hausdorff dimension zero.
It has come to our attention that our construction has some similarities to that of [2], but our series expansion of real numbers is inspired by the binary expansion. Erdős and Volkmann [2] noted that the question of the Hausdorff dimension of subrings and subfields remained open. We feel that it is of interest to see how Lemma 2 is crucially used to show that we get a subring of $\mathbb{R}$ by using numbers which are represented by special series expansions. Also, we are able to see relatively easily that the set AA has Hausdorff dimension zero.
We have been informed of the following construction. If AA is a compact Cantor set such that for every $n\in\mathbb{N}$ the set $A^{n}$ has Hausdorff dimension zero, then it can be shown that the subring of $\mathbb{R}$ which is generated by AA has Hausdorff dimension zero. However, we are unable to find a reference for the construction of such a set. Our construction is very explicit using series expansions and we use neither the Continuum Hypothesis nor the Axiom of choice.